OpenStudy (anonymous):

answer and i will give you a metal part 1 What are the steps to find the inverse of a function equation? For example, what is the inverse of f(x) = 2x -7? part 2 Explain how do you find the inverse of a point. For example, what is the inverse of (0,2) and (-6,0).

OpenStudy (anonymous):

function: f(x)=2x-7 y=2x-7 to make it inverse, switch around the x and y. so x=2y-7 x+7=2y y=1/2x+7/2. to find the inverse of a point, also switch around x and y. so (0,2) would become (2,0) and (-6,0) would become (0,-6).

OpenStudy (anonymous):

can you tell me what the steps are

OpenStudy (anonymous):

all you do is switch the y and the x around it's not really math.. :P do you get it?

OpenStudy (anonymous):

yeah thank you